# Synthetic Query Generation for Privacy-Preserving Deep Retrieval Systems using Differentially Private Language Models

### v2

@@ Abstract
We address the challenge of ensuring differential privacy (DP) guarantees in training deep retrieval systems. Training these systems often involves the use of contrastive-style losses, which are typically non-per-example decomposable, making them difficult to directly DP-train with since common techniques require per-example gradient. To address this issue, we propose an approach that prioritizes ensuring query privacy prior to training a deep retrieval system. Our method employs DP language models (LMs) to generate private synthetic queries representative of the original data. These synthetic queries can be used in downstream retrieval system training without compromising privacy. Our approach demonstrates a significant enhancement in retrieval quality compared to direct DP-training, all while maintaining query-level privacy guarantees. This work highlights the potential of harnessing LMs to overcome limitations in standard DP-training methods.
@@
Alex Kurakin Google Research kurakin@google.com                        Matthew Jagielski Google Research jagielski@google.com                        Milad Nasr Google Research srxzr@google.com
@@ 1 Introduction
The standard approach to ensure the privacy of training data in large-scale machine learning models is to directly introduce differential privacy (DP) (Dwork et al., 2014) guarantees during training (Abadi et al., 2016; Ponomareva et al., 2023). These DP-training strategies provide guarantees by limiting the impact each individual data instance has on the overall model. However, some models contain design elements that inherently hinder the ability to limit per-example contributions, and thus are more difficult to directly DP-train. These include models with components that calculate batch statistics like batch normalization layers and losses that cannot be decomposed into per-example losses like pairwise and contrastive-style losses (Huai et al., 2020; Xue et al., 2021; Ponomareva et al., 2023). Hence, additional considerations are often needed to adapt DP-training to such models for achieving an adequate privacy-accuracy tradeoff.
For non-per-example decomposable losses in particular, the injected noise needed to achieve DP guarantees typically scales with the number of examples that appear in example-level loss computations, which may result in excessive model utility degradation. For instance, deep retrieval systems often employ non-per-example decomposable contrastive-style losses to train semantic neural representations of user queries and candidate items to facilitate efficient vector-based retrieval strategies. Deep retrieval systems have been widely adopted in many online services from search to advertising (Covington et al., 2016; Huang et al., 2020). Although these personalized systems enhance user experience, it is crucial that this improvement does not compromise user privacy. The ubiquity of these systems underscores the importance of developing general-purpose approaches to ensure privacy guarantees without significantly hindering their utility—an area that has seen limited research efforts.
@@ 2 Related Work
Several recent studies (Yue et al., 2022; Mattern et al., 2022; Mireshghallah et al., 2022; Putta et al., 2023) have investigated the utility of private synthetic data from DP finetuned LMs in downstream tasks with pointwise losses, including text classification and semantic parsing. These works find that downstream models trained with private synthetic data outperform directly DP-trained models under the same privacy budget, and non-private synthetic data generation can even improve performance under no privacy constraints. The reason is that DP synthetic data benefits from the injection of additional public information from the pretrained LMs. Our work contributes another exploration of the advantages of private synthetic data generation for downstream training under a different learning paradigm with a non-per-example decomposable loss.
@@ 3.1 Deep Retrieval
Deep retrieval systems have emerged as highly effective and scalable information retrieval systems to find candidate items based on their semantic relevance to a specific query (Huang et al., 2020; Ni et al., 2021). These systems typically consist of two neural encoders capable of generating rich, dense representations of queries and items (see Figure 2), which enable approximate nearest neighbor search methods to efficiently retrieve semantically relevant items to a given query. Deep retrieval systems are typically trained on contrastive-style losses that make use of two types of data examples: positive examples and negative examples. The positive examples help train the encoders into pulling relevant query-item pair embeddings close together in the embedding space, while negative examples help in preventing embedding space collapse. A popular choice for the loss function in deep retrieval is the in-batch softmax loss, which makes memory-efficient use of items already loaded in a mini-batch as randomly sampled soft negatives (Gillick et al., 2019; Karpukhin et al., 2020; Qu et al., 2020). In particular, given a training batch of query-item pairs $\{(q_{i},d_{i})\}_{i\in\mathcal{B}}$, each $d_{i}$ is the positive item document for query $q_{i}$, and all other item documents $\{d_{j}\}_{j\neq i}$ within the batch are treated as the negatives. The in-batch softmax loss for each sample in the batch is
@@
Figure 2: Illustration of deep retrieval dual encoder model. Dashed lines connecting the encoders represent a shared encoder.
@@ 3.3 Differential Privacy
This definition captures a privacy guarantee based on the indistinguishability of the presence of a single data point in a dataset. The $\epsilon$ and $\delta$ parameter control the strength of this privacy guarantee, where smaller values correspond to stronger guarantees. A useful property of DP that is crucial to our approach is its post-processing property (Dwork et al., 2014) which states that for any deterministic or randomized function $f$ defined over the range of the mechanism $\mathcal{M}$, if $\mathcal{M}$ satisfies $(\epsilon,\delta)$-DP, so does the composition $f\circ\mathcal{M}$. The post-processing property ensures that arbitrary computations on the output of a DP mechanism do not incur any additional privacy loss.
@@ 3.3.2 Limitations of Directly Differentially Private Training Retrieval Systems
Our work was primarily motivated by the fact that DP-SGD is not immediately compatible with the in-batch softmax loss used to train dual encoders. The primary reason is that per-example gradients of this loss depend on not just the example in consideration but also all other examples in the batch. This results in increased sensitivity and thus greater noise additions that tend to worsen utility.
@@
Lastly, a systems-level issue of DP-training on the in-batch softmax loss is that in order to take advantage of vectorization and parallelization strategies for computing per-example gradients more quickly (Subramani et al., 2021), each query-item example in a batch must be duplicated to be contained in every de facto example in the batch, leading to a quadratic increase in memory requirements. Given fixed memory resources, this necessitates significantly smaller batch sizes, which has an additional deleterious effect beyond gradient clipping and noising since effective representation learning under the in-batch softmax loss highly depends on the amount and diversity of in-batch examples. Our approach of training with private synthetic queries precludes the above limitations when training a downstream dual encoder.
@@ 4 Approach
Then, the DP fine-tuned LM is capable of generating synthetic queries that are representative of the real queries and relevant to the items. For each item document $d$, we generate a matching synthetic query $\tilde{q}$ by providing the input “generate_query: $d$” to the model. A synthetic training dataset is then constructed to be the set of original documents matched with their corresponding synthetic queries.
@@ 5.2.2 Data Synthesis
Source
Text
Document
The main cast of the show: Mickey Mouse, Minnie Mouse, Donald Duck, Daisy Duck, Goofy, and Pluto star in the series, which focuses on interacting with the viewer to stimulate problem solving.
Original Query
characters from the Mickey Mouse clubhouse show
T5-Base
$\epsilon=\infty$
the Mickey Mouse show cast
$\epsilon=16$
what is the most important characters in the series
$\epsilon=8$
what is in this series?
$\epsilon=3$
what is in this code for dfr1
T5-XL
$\epsilon=3$
issue with show by mickey mouse
T5-Large
$\epsilon=3$
what is the most animated characters on disney cartoon show
T5-Small
$\epsilon=3$
what is isn’t a character will do a story
@@ 6.1.1 MSMARCO Evaluation
Table 3 shows the evaluation results on the MSMARCO test set for deep retrieval models regularly trained on the synthetically generated data under varying generative model configurations and varying privacy levels. For our benchmark comparison, in Table 4 we display the evaluation results on the same test set of deep retrieval models DP-trained on the original data under varying privacy levels. In the top rows of both tables, we provide a baseline reference evaluation of a dual encoder model regularly trained on the original data without any DP guarantees (i.e., $\epsilon=\infty$).
@@ 7 Conclusion
There are a few limitations to our approach. Firstly, while we observed that larger LMs generate higher quality synthetic queries, it is worth noting that training such large models may too computationally expensive. Exploring more parameter-efficient finetuning methods tailored to DP-training could mitigate the computational burden associated with training such larger models. Secondly, it is necessary for the publicly pretrained LM utilized for generating synthetic queries to not have been pretrained on the original queries we aim to privatize. This imposes a constraint on the choice of LMs suitable for generating private synthetic queries. Next, as observed in the synthetic example in Table 2, the DP-finetuned LMS can sometimes generate incoherent queries, which can limit the relevance and interpretability of the data. Lastly, it is essential to recognize that our approach exclusively ensures query-level privacy. For achieving more general example-level privacy, it is necessary to resort to more conventional DP training methods.
@@
Language models have been demonstrated to implicitly memorize sensitive user information in the training data Carlini et al. (2021, 2023). Therefore, the introduction of these high-capacity models into retrieval and recommendation systems poses serious privacy risks. Data privacy is a crucial consideration in the responsible development of these personalized systems. Our work directly offers potential solutions to address these issues. However, it is important to acknowledge limitations on privacy guarantees our approach to prevent undesired privacy risks.

### v3

@@ Abstract
We address the challenge of ensuring differential privacy (DP) guarantees in training deep retrieval systems. Training these systems often involves the use of contrastive-style losses, which are typically non-per-example decomposable, making them difficult to directly DP-train with since common techniques require per-example gradients. To address this issue, we propose an approach that prioritizes ensuring query privacy prior to training a deep retrieval system. Our method employs DP language models (LMs) to generate private synthetic queries representative of the original data. These synthetic queries can be used in downstream retrieval system training without compromising privacy. Our approach demonstrates a significant enhancement in retrieval quality compared to direct DP-training, all while maintaining query-level privacy guarantees. This work highlights the potential of harnessing LMs to overcome limitations in standard DP-training methods.
@@
Alex Kurakin Google DeepMind kurakin@google.com                        Matthew Jagielski Google DeepMind jagielski@google.com                        Milad Nasr Google DeepMind srxzr@google.com
@@ 1 Introduction
Deep retrieval systems have been widely adopted in many online services, from search to advertising, to match user queries to relevant recommendations (Covington et al., 2016; Huang et al., 2020). In many applications, candidate items for retrieval are often publicly available non-personal information in the sense that they do not contain any specific information related to any single user (e.g., articles, products, movies, ads). However, the input queries to retrieval systems can often contain user personal information. Therefore, training deep retrieval systems on user data may enhance user experience through timely relevance, but it may also unintentionally compromise user privacy since neural network models have been demonstrated to implicitly memorize and leak sensitive user information in the training data (Carlini et al., 2019). This raises privacy sensitivities around each stage of data collection, training, inference, and hosting these systems. In this work, we seek to address the problem of ensuring user query privacy guarantees in deep retrieval systems without significantly hindering their utility.
The standard approach to ensure the privacy of training data in many large-scale machine learning models is to directly introduce differential privacy (DP) (Dwork et al., 2014) guarantees during training (Abadi et al., 2016; Ponomareva et al., 2023). These DP-training strategies provide guarantees by limiting the impact each individual data instance has on the overall model. However, some models contain design elements that inherently hinder the ability to limit per-example contributions, and thus are more difficult to directly DP-train. These include models with components that calculate batch statistics such as batch normalization layers (Ponomareva et al., 2023) and models with losses that cannot be decomposed into per-example losses such as pairwise and contrastive-style losses (Huai et al., 2020; Xue et al., 2021).
This limits the application of DP-training on deep retrieval systems since these systems typically use non-per-example decomposable contrastive-style losses to train semantic neural representations of user queries and candidate items in order to facilitate efficient vector-based retrieval strategies. The injected noise needed to achieve DP guarantees for these losses can scale with the number of candidate items that appear in the example-level loss computations, which can result in excessive retrieval quality degradation. Hence, additional considerations are often needed to adapt DP-training to deep retrieval models for achieving an adequate privacy-performance tradeoff.
Figure 1: Illustration of approach.
@@ 2 Related Work
Several recent studies (Yue et al., 2022; Mattern et al., 2022; Mireshghallah et al., 2022; Putta et al., 2022) have investigated the utility of private synthetic data from DP finetuned LMs in downstream tasks with pointwise losses, including text classification and semantic parsing. These works find that downstream models trained with private synthetic data outperform directly DP-trained models under the same privacy budget, and non-private synthetic data generation can even improve performance under no privacy constraints. The reason is that DP synthetic data benefits from the injection of additional public information from the pretrained LMs. Our work contributes another exploration of the advantages of private synthetic data generation for downstream training under a different learning paradigm with a non-per-example decomposable loss. In particular, our motivation is to achieve query privacy DP guarantees in deep retrieval systems with high utility.
@@ 3.1 Deep Retrieval
Deep retrieval systems have emerged as highly effective and scalable information retrieval systems to find candidate items based on their semantic relevance to a specific query (Huang et al., 2020; Ni et al., 2021). These systems typically consist of two neural encoders capable of generating rich, dense representations of queries and items (see Figure 2), which enable efficient approximate nearest neighbor search methods (Guo et al., 2020) to retrieve semantically relevant items to a given query. Deep retrieval systems are typically trained on contrastive-style losses that make use of two types of data examples: positive examples and negative examples. The positive examples help train the encoders into pulling relevant query-item pair embeddings close together in the embedding space, while negative examples help in preventing embedding space collapse. A popular choice for the loss function in deep retrieval is the in-batch softmax loss, which makes memory-efficient use of items already loaded in a mini-batch as randomly sampled soft negatives (Gillick et al., 2019; Karpukhin et al., 2020; Qu et al., 2020). In particular, given a training batch of query-item pairs $\{(q_{i},d_{i})\}_{i\in\mathcal{B}}$, each $d_{i}$ is the positive item document for query $q_{i}$, and all other item documents $\{d_{j}\}_{j\neq i}$ within the batch are treated as the negatives. The in-batch softmax loss for each sample in the batch is
@@
Figure 2: Illustration of deep retrieval dual encoder model. Dashed lines connecting the query and document encoders represent the possibility of sharing the same encoder.
3.1.1 Privacy Risks of Deep Retrieval
The neural encoders in deep retrieval systems are high-capacity models known to implicitly memorize sensitive information present in the training data (Carlini et al., 2019). Such sensitive data can subsequently be extracted from these trained models (Carlini et al., 2021; Lehman et al., 2021). Moreover, retrieval-augmented text generation systems, which utilize deep retrieval systems to aid text generation, have been demonstrated to be more susceptible to leaking private information from their private datastore compared to the language models trained on the private data (Huang et al., 2023; Zeng et al., 2024). This underscores the increase in privacy risks associated with deep retrieval systems.
@@ 3.3 Differential Privacy
Note that for capturing query-level differential privacy in retrieval datasets under this definition, neighboring datasets are datasets that differ by exactly one query. This definition captures a privacy guarantee based on the indistinguishability of the presence of a single data point in a dataset. The $\epsilon$ and $\delta$ parameter control the strength of this privacy guarantee, where smaller values correspond to stronger guarantees. A useful property of DP that is crucial to our approach is its post-processing property (Dwork et al., 2014) which states that for any deterministic or randomized function $f$ defined over the range of the mechanism $\mathcal{M}$, if $\mathcal{M}$ satisfies $(\epsilon,\delta)$-DP, so does the composition $f\circ\mathcal{M}$. The post-processing property ensures that arbitrary computations on the output of a DP mechanism do not incur any additional privacy loss.
@@ 3.3.2 Limitations of Directly Differentially Private Training Retrieval Systems
Our work was primarily motivated by the fact that DP-SGD is not immediately compatible with the in-batch softmax loss used to train dual encoders. The primary reason is that per-example gradients of this loss depend on not just the example in consideration but also all other examples in the batch. Therefore, a single example can influence multiple per-example gradient computations which immediately implies an increased sensitivity of the gradient that scales with the batch size. In DP-training, higher sensitivity means that more noise needs to be added to the gradient updates during training to achieve the same level of privacy guarantees, which leads to worse utility.
@@
Lastly, a systems-level issue of DP-training on the in-batch softmax loss is that in order to take advantage of vectorization and parallelization strategies for computing per-example gradients more quickly (Subramani et al., 2021), each query-item example in a batch must be duplicated to be contained in every de facto example in the batch, leading to a quadratic increase in memory requirements. Given fixed memory resources, this necessitates significantly smaller batch sizes, which has an additional deleterious effect beyond gradient clipping and noising since effective representation learning under the in-batch softmax loss highly depends on the amount and diversity of in-batch examples. Our approach of training with private synthetic queries precludes the above limitations when training a downstream dual encoder deep retrieval model.
@@ 4 Approach
Then, the DP fine-tuned LM is capable of generating synthetic queries that are representative of the real queries and relevant to the items. For each item document $d$, we generate a matching synthetic query $\tilde{q}$ by providing the input “generate_query: $d$” to the model. This method allows for generating multiple synthetic queries from each document. A synthetic training dataset is then constructed to be the set of original documents matched with their corresponding synthetic queries.
@@ 5.2.2 Data Synthesis
Source Text
Document The main cast of the show: Mickey Mouse, Minnie Mouse, Donald Duck, Daisy Duck, Goofy, and Pluto star in the series, which focuses on interacting with the viewer to stimulate problem solving.
Original Query characters from the Mickey Mouse clubhouse show
T5-Base $\epsilon=\infty$ the Mickey Mouse show cast
$\epsilon=16$ what is the most important characters in the series
$\epsilon=8$ what is in this series?
$\epsilon=3$ what is in this code for dfr1
T5-XL $\epsilon=3$ issue with show by mickey mouse
T5-Large $\epsilon=3$ what is the most animated characters on disney cartoon show
T5-Small $\epsilon=3$ what is isn’t a character will do a story
@@ 6.1 Evaluation on Retrieval Tasks
6.1.1 Search & Retrieval Procedure
Evaluation of the dual encoder retrieval models requires a query-document nearest neighbor search implementation for the inference stage. For our experiments, we used the Scalable Nearest Neighbors (ScaNN) library, an open-source library that provides a fast and scalable approximate nearest neighbor search procedure (Guo et al., 2020). The procedure was executed using ScaNN’s brute force scoring and the inner product distance settings.
@@ 6.1.2 MSMARCO Evaluation
Table 3 shows the evaluation results on the MSMARCO test set for deep retrieval models regularly trained on the synthetically generated data under varying generative model configurations and varying privacy levels. For our benchmark comparison, in Table 4 we display the evaluation results on for the deep retrieval models directly DP-trained on the original data under varying privacy levels. In the top rows of both tables, we also provide another baseline reference evaluation of a dual encoder model regularly finetuned on the original data without any DP guarantees (i.e., $\epsilon=\infty$).
@@ 7 Conclusion
There are a few limitations to our approach. Firstly, while we observed that larger LMs generate higher quality synthetic queries, it is worth noting that training such large models may too computationally expensive. Exploring more parameter-efficient finetuning methods tailored to DP-training could mitigate the computational burden associated with training such larger models. Secondly, it is necessary for the publicly pretrained LM utilized for generating synthetic queries to not have been pretrained on the original queries we aim to privatize. This imposes a constraint on the choice of pretrained LMs suitable for generating private synthetic queries. Next, as observed in the synthetic example in Table 2, the DP-finetuned LMs can sometimes generate incoherent queries, which can limit the relevance and interpretability of the data. Lastly, it is essential to recognize that our approach exclusively ensures query-level privacy. For achieving more general example-level privacy, it may be necessary to resort to other approaches or more conventional DP-training methods.
@@
Data privacy is a crucial consideration in the responsible development of personalized machine learning systems. Our work directly offers potential solutions to address issues of data privacy in deep retrieval systems. However, it is important to acknowledge the above limitations on privacy guarantees of our approach to prevent undesired privacy risks.
Acknowledgements
We thank Rishabh Bansal, Manoj Reddy, Andreas Terzis, Sergei Vassilvitskii, Abhradeep Guha Thakurta, Shuang Song, Arthur Asuncion, and Heather Yoon for the helpful discussions. We also thank Jianmo Ni for their assistance in setting up the retrieval training pipeline. Finally, we appreciate the support and encouragement of YouTube Ads leadership Shobha Diwakar, Marija Mikic, and Ashish Gupta throughout this work.